OpenStudy (anonymous):

Any help would be much appreciated..... x/5+4=5/3

OpenStudy (bahrom7893):

multiply everything by 5: x + 20 = 25/3 multiply everything by3: 3x + 60 = 25

OpenStudy (bahrom7893):

subtract 60 from both sides: 3x + 60 - 60 = 25 - 60 3x = -35

OpenStudy (bahrom7893):

Divide both sides by 3: x = -35/3

OpenStudy (bahrom7893):

please click on become a fan if I helped, THANKS!=)